# Generalized local jocabians and commutative group stacks

### v2

@@ 1.3. Commutative algebraic group stacks
from affine $k$-schemes to connective complexes of abelian groups, satisfying the usual $fppf$ hyper-descent condition. Any object $E\in\mathbf{Ab}_{k}$ possesses an underlying derived stack $|E|\in\mathbf{St}_{k}$ obtained by the Dold-Kan construction, explicitly defined by
$|E|(S):=\mathbf{Map}_{\mathbf{dg}_{\mathbb{Z}}}(\mathbb{Z},E(S))$
for any $S\in\mathbf{Aff}_{k}$. This defines a forgetful $\infty$-functor $\mathbf{Ab}_{k}\to\mathbf{dSt}_{k}$, which admits a left adjoint sending $F$ to $\mathbb{Z}\otimes F$. Recall that for $E\in\mathbf{AbSt}_{k}$ we denote by $\pi_{i}(E)$ the cohomology sheaves $H^{-i}(E)$, which are also the homotopy sheaves $\pi_{i}(|E|)$ of the underlying stack $|E|$ pointed at 00.
@@
A commutative algebraic group stack is an object $E\in\mathbf{AbSt}_{k}$ such that underlying stack $|E|\in\mathbf{St}_{k}$ is an Artin $n$-stack (for some $n$), locally of finite presentation and strongly quasi-separated over $k$. They form, by definition, a full sub-$\infty$-category
@@
Remind that strongly quasi-separated means that the diagonal map $E\to E\times E$ is strongly quasi-compact in the sense of [TV08]. This means that all the higher diagonal maps $X\to X^{S^{n}}$ are quasi-compact for $n\geq 0$. Equivalently, the inertia stack $IE\to E$ is required to be strongly quasi-compact over $E$. As explained in [Toe05, Cor. 2.9] an Artin stack $E$ locally of finite presentation which is strongly quasi-separated is such that for all field value point $s\in E(K)$, all its homotopy sheaves $\pi_{i}(E,s)$ are quasi-compact and quasi-separated algebraic spaces over $K$, and thus are representable by group schemes thanks to [Art69, Lem. 4.2]. As we are working over a field, the converse also holds.
@@
An object $E\in\mathbf{AbSt}_{k}$ is a commutative algebraic group stack if and only if each homotopy sheaf $\pi_{i}(E)\in Ab_{k}$ is representable by a group scheme locally of finite type over $k$, which is moreover quasi-projective for $i>0$ and vanish for all but a finite number of indices $i$.
Proof. Assume first that $E$ is a commutative algebraic group stack. In the definition 1.3. We have to prove that all the homotopy sheaves $\pi_{i}(E)$ are representable. We consider the inertia stack $IE$ of $E$. We already saw that $\pi_{i}(E)$ are group schemes for $i>0$ by means of [Toe05, Cor. 2.9], it remains to treat the case $i=0$. As $E$ is an abelian group object we have splitting $IE\simeq E\times\Omega_{0}E$, where $\Omega_{0}E$ is the pointed loop stack at 00. As $k$ is a field, we thus have that $IE\to E$ is a fppf cover which is strongly quasi-compact. We can thus use [Toe05, Prop. 2.4], which remains valid only assuming the diagonal to be strongly quasi-compact, and deduce that $E$ is a gerbe. In particular, $\pi_{0}(E)$ is representable by an algebraic space locally of finite type over $k$. This is a group object in algebraic space, and is moreover quasi-separated by assumptions on $E$, and by [Art69, Lem. 4.2] is representable by a group scheme locally of finite type.
Conversely, assume that $E$ satisfies the conditions of the proposition. We will prove by induction on $i$ that the $i$-th homotopical truncation $\tau_{\leq i}E$ is an Artin $i$-stack locally of finite presentation and strongly quasi-separated over $k$. For $i=0$ this is obvious, as $\pi_{0}(E)$ is a scheme locally of finite type by assumption. Consider $E_{\leq i}\to E_{\leq i-1}$, which fits into a cartesian square
[…]
By assumption, $\pi_{i}(E)$ is a group scheme of finite type over $k$ for $i>0$, and thus $K(\pi_{i}(E),i+1)$ is an Artin $(i+1)$-stack locally of finite presentation and strongly quasi-separated $k$. By induction, this implies that $E_{\leq i}$ is an Artin $i$-stack locally of finite presentation and strongly quasi-separated. $\Box$
@@
The full sub-$\infty$-category $\mathbf{AbSt}^{alg}_{k}\subset\mathbf{AbSt}_{k}$ is a is stable by finite limits, retracts and the homotopical truncation $\infty$-functors $\tau_{\leq i}$.
@@
We warn the reader that if $E$ is a commutative algebraic group stack then $E[1]=BE$ is in general not a commutative algebraic group stack anymore. It is so only if $\pi_{0}(E)$ is moreover assumed to be of finite type.
@@
A commutative locally algebraic group stack is an object $E\in\mathbf{AbSt}_{k}$ for all $i$ the sheaf $\pi_{i}(E)$ is representable by an algebraic space locally of finite type. The $\infty$-category of commutative locally algebraic group stacks is denoted by $\mathbf{AbSt}^{lalg}_{k}\subset\mathbf{AbSt}_{k}$.
As opposed to algebraic group stacks, the locally algebraic stacks are stable by suspension: if $E\in\mathbf{AbSt}_{k}^{lalg}$, then $BE$, computed in $\mathbf{AbSt}_{k}$, remains in $\mathbf{AbSt}_{k}^{lalg}$. More generally, $\mathbf{AbSt}_{k}^{lalg}$ is stable by finite limits and colimits, as this can be easily seen using the long exact sequence in homotopy.
@@ 2. Algebraic homology
defined by $E\mapsto\mathbf{Map}(X,E)$. This $\infty$-functor will be denoted by $E\mapsto\mathbb{H}(X,E)$. By definition it obviously commutes with finite limits, and thus is pro-representable by an object
@@
By definition we have in particular that for any commutative group scheme $G$, separated and of finite type over $k$
@@
The notion above of algebraic homology is very close to that of flat homology of [Mil76], but is different. The major difference is that these two objects live in two different, non-equivalent, $\infty$-categories. Flat homology lives in the $\infty$-category of complexes of commutative pro-algebraic groups. This has a canonical $\infty$-functor to $\mathbf{AbSt}_{k}$, by considering the stack pro-represented by a complex of pro-alegebraic groups, but this $\infty$-functor is known not to be fully faithful. This is due to the existence of two different, non-isomorphic, definitions of ext-groups of commutative algebraic groups: Yoneda ext-groups inside algebraic groups and ext-groups of the corresponding fppf-sheaves. It is known that these two notions are different (see for instance [Bre69]).
@@
Functoriality. The construction $X\mapsto\mathbb{H}^{alg}(X)$ is functorial in $X$, and defines an $\infty$-functor from Artin $k$-stack to pro-object in derived commutative algebraic groups. Moreover, the identity of $\mathbb{H}^{alg}(X)$ defines a canonical class $\alpha_{X}\in\mathbb{H}(X,\mathbb{H}^{alg}(X))$, and thus a canonical morphism of pro-stacks
@@
Homology groups. The pro-object $\mathbb{H}^{alg}(X)$ possesses pro-homology groups $\mathbb{H}_{i}^{alg}(X):=\pi_{i}(\mathbb{H}^{alg}(X))$, which are commutative pro-algebraic groups over $k$. By definition our notion of commutative algebraic group stacks $\mathbb{H}_{0}^{alg}(X)$ is a pro (separated) group scheme locally of finite presentation, and all the $\mathbb{H}_{i}^{alg}(X)$ are group schemes separated and of finite type for $i>0$.
@@
$0\to\mathbb{H}_{0}^{alg}(X)^{(0)}\to\mathbb{H}_{0}^{alg}(X)\to\mathbb{Z}\to 0$
We can consider the universal morphism $X\to\mathbb{H}^{alg}(X)$, and check that the composite morphism $X\to\mathbb{H}_{0}^{alg}(X)\to\mathbb{Z}$ is the constant function $1$ on $X$ as soon as $X$ is geometrically connected. It thus factors through a canonical morphism
@@
where $\mathbb{H}_{0}^{alg}(X)^{(1)}$ is the fiber at $1$ of the short exact sequence above, considered naturally as a torsor over the group $\mathbb{H}_{0}^{alg}(X)^{(0)}$. This morphism is a generalization of the usual Albanese morphism to the Albanese variety. When $X$ is smooth and proper, $\mathbb{H}_{0}^{alg}(X)^{(1)}$ can be identified with the usual Albanese variety, and $alb_{X}$ with the usual the Albanese map, as these two objects satisfy the same universal property.
In general, the universal morphism $X\to\mathbb{H}^{alg}(X)$, composed with the projection $\mathbb{H}^{alg}(X)\to\mathbb{Z}$ is the constant map $X\to\mathbb{Z}$ equals to $1$. Therefore, we have a canonical morphism $X\to\mathbb{H}^{alg}(X)^{(1)}$ where $\mathbb{H}^{alg}(X)^{(1)}$ is the fiber at $1$ of $\mathbb{H}^{alg}(X)\to\mathbb{Z}$. This morphism is a higher cohomological generalization of the Albanese morphism as being the universal map towards a torsor over an object in $Pro(\mathbf{AbSt}_{k}^{alg})$.
Relation to the Albanese stack. When $X$ is an Artin stack of finite type over $X$, the first Postnikov truncation $\mathbb{H}_{\leq 1}^{alg}(X)$ is an abelian group object in algebraic pro-Artin $1$-stacks. This object is closely related $D(Pic(X))$, the dual of the Picard stack of $X$ (relative to $k$) studied in [Bro21]. Indeed, $Pic(X):=\underline{\mathbf{Map}}(X,B\mathbb{G}_{m})$ is the internal $Hom$ stack from $X$ to $B\mathbb{G}_{m}$, and we thus have a canonical evaluation morphism $ev:X\times Pic(X)\to B\mathbb{G}_{m}$ which defines a canonical morphism
$X\to\underline{\mathbf{Map}}(Pic(X),B\mathbb{G}_{m})=D(Pic(X)).$
When $X$ is proper and the dual $D(Pic(X))$ is known to be a representable group stack and sits in a fibration sequence of commutative group stacks
$0\to D(\pi_{0}(Pic(X)))\to D(Pic(X))\to\mathbb{Z}\to 0$
The choice of a global point $x\in X(k)$ defines a splitting of this sequence, and thus the evaluation map $X\to D(Pic(X))$ can be composed to a morphism $X\to D(H^{0}(Pic(X))$. By the universal property of algebraic homology we get a canonical morphism of pro-objects in commutative algebraic $1$-stacks
$\mathbb{H}_{\leq 1}^{alg}(X)^{(0)}\to D(H^{0}(Pic(X))).$
In absence of global points we get a morphism $\mathbb{H}_{\leq 1}^{alg}(X)^{(1)}\to D(H^{0}(Pic(X)))$ where $\mathbb{H}_{\leq 1}^{alg}(X)^{(1)}$ is the fiber at $1$ of the canonical fibration sequence
@@
as described above. This map is an isomorphism when $X$ satisfies some extra conditions expressed in [Bro21] (e.g. when $X$ is smooth and proper over $k$), as both objects $\mathbb{H}_{\leq 1}^{alg}(X)^{(1)}$ and $D(H^{0}(Pic(X)))$ then satisfy the same universal property. This provides a direct relation between algebraic homology and the Albanese morphism constructed in [Bro21].
@@
We claim that this morphism is in fact an equivalence of pro-object. Indeed, for any $G\in AbSt_{k}^{alg}$, we have

### v3

@@ 1.3. Stacks
from the category of affine $k$-schemes to the category $\mathbf{dg}^{c}_{\mathbb{Z}}$ of connective complexes of abelian groups, satisfying the previously mentioned fppffppf hyper-descent condition. Any object $E\in\mathbf{Ab}_{k}$ possesses an underlying stack $|E|\in\mathbf{St}_{k}$ obtained by the Dold-Kan construction, explicitly defined by
$|E|(X):=\mathbf{Map}_{\mathbf{dg}^{c}_{\mathbb{Z}}}(\mathbb{Z},E(X))$
for any $X\in\mathbf{Aff}_{k}$, where $\mathbf{Map}_{\mathbf{dg}^{c}_{\mathbb{Z}}}$ denotes the mapping spaces of the $\infty$-category $\mathbf{dg}^{c}_{\mathbb{Z}}$. This defines a forgetful $\infty$-functor $\mathbf{Ab}_{k}\to\mathbf{dSt}_{k}$, which admits a left adjoint sending $F$ to $\mathbb{Z}\otimes F$. For $E\in\mathbf{AbSt}_{k}$ we denote by $\pi_{i}(E)$ the cohomology sheaves $H^{-i}(E)$, which are canonically identified with the homotopy sheaves $\pi_{i}(|E|)$ of the underlying stack $|E|$ pointed at global point 0. Finally, for $E\in\mathbf{AbSt}_{k}$, the cartesian square coming from the Postnikov decompositions admit canonical promotion to cartesian square in $\mathbf{AbSt}_{k}$
${\tau_{\leq n}(E)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\tau_{\leq n-1}(E)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{k_{n}}\textstyle{\bullet\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{n}(E)[n+1]}$
(note that the underlying stack of $\pi_{n}(E)[n+1]$ is $K(\pi_{n}(E),n+1)$).
1.4. Commutative algebraic group stacks
We remind from [31, §2.2.1] the notion of Artin $n$-stack.
@@ 1.4. Commutative algebraic group stacks
A commutative algebraic group stack is an object $E\in\mathbf{AbSt}_{k}$ such that the underlying stack $|E|\in\mathbf{St}_{k}$ is an Artin $n$-stack (for some $n$), locally of finite presentation, strongly quasi-separated over $k$ and which can be covered, in the sense of the fppf topology, by a countable number of affine $k$-schemes. They form, by definition, a full sub-$\infty$-category
@@
We recall that strongly quasi-separated means that the diagonal map $E\to E\times E$ is strongly quasi-compact in the sense of [31]. This means that all the higher diagonal maps $X\to X^{S^{n}}$ are quasi-compact for $n\geq 0$. Equivalently, the inertia stack $IE\to E$ is required to be strongly quasi-compact over $E$. The countability assumption is here to ensure that commutative algebraic group stacks form an essentially small $\infty$-category.
As explained in [29, Cor. 2.9] an Artin stack $E$ locally of finite presentation which is strongly quasi-separated is such that for all field valued point $s\in E(K)$, all its homotopy sheaves $\pi_{i}(E,s)$ are quasi-compact and quasi-separated algebraic spaces of finite ype over $K$, and thus are representable by group schemes thanks to [1, Lem. 4.2].
As we are working over a field, the converse also holds.
@@
An object $E\in\mathbf{AbSt}_{k}$ is a commutative algebraic group stack if and only if each homotopy sheaf $\pi_{i}(E)\in Ab_{k}$ is representable by a group scheme locally of finite type over $k$, which is moreover quasi-projective for $i>0$ and vanishes for all but a finite number of indices $i$.
Proof. Assume first that $E$ is a commutative algebraic group stack. In the definition 1.3. We have to prove that all the homotopy sheaves $\pi_{i}(E)$ are representable. We consider the inertia stack $IE$ of $E$. We already saw that $\pi_{i}(E)$ are group schemes for $i>0$ by means of [29, Cor. 2.9]111the result of [29, Cor. 2.9] assumes the quasi-compactness of $E$, but the result remains valid in our situation as any point of $E$ with value in an affine scheme will factor through a quasi-compact open substack $E^{\prime}\subset E$., it remains to treat the case $i=0$. As $E$ is an abelian group object we have splitting $IE\simeq E\times\Omega_{0}E$, where $\Omega_{0}E$ is the pointed loop stack at 0. As $k$ is a field, we thus have that $IE\to E$ is a fppf cover which is strongly quasi-compact. We can thus use [29, Prop. 2.4], which remains valid only assuming the diagonal to be strongly quasi-compact, and deduce that $E$ is a gerbe. In particular, $\pi_{0}(E)$ is representable by an algebraic space locally of finite type over $k$. This is a group object in algebraic space, and is moreover quasi-separated by assumptions on $E$, and by [1, Lem. 4.2] is representable by a group scheme locally of finite type.
Conversely, assume that $E$ satisfies the conditions of the proposition. We will prove by induction on $i$ that the $i$-th homotopical truncation $\tau_{\leq i}E$ is an Artin $i$-stack locally of finite presentation and strongly quasi-separated over $k$. For $i=0$ this is obvious, as $\tau_{\leq 0}E\simeq\pi_{0}(E)$ is a scheme locally of finite type by assumption. Consider $\tau_{\leq i}E\to\tau_{\leq i-1}E$, which fits into a cartesian square
${\tau_{\leq i}E\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\tau_{\leq i-1}E\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{i}(E)[i+1].}$
By assumption, $\pi_{i}(E)$ is a group scheme of finite type over $k$ for $i>0$, and thus $K(\pi_{i}(E),i+1)$ is an Artin $(i+1)$-stack locally of finite presentation and strongly quasi-separated over $k$. By induction, this implies that $\tau_{\leq i}E$ is an Artin $i$-stack locally of finite presentation and strongly quasi-separated. $\Box$
@@
The full sub-$\infty$-category $\mathbf{AbSt}^{alg}_{k}\subset\mathbf{AbSt}_{k}$ is is stable by finite limits, retracts and the homotopical truncation $\infty$-functors $\tau_{\leq i}$.
@@
We caution the reader that if $E$ is a commutative algebraic group stack then $E[1]=BE$ is generally not a commutative algebraic group stack anymore. It is so only if $\pi_{0}(E)$ is moreover assumed to be of finite type.
@@
A commutative locally algebraic group stack is an object $E\in\mathbf{AbSt}_{k}$ such that for all $i$ the sheaf $\pi_{i}(E)$ is representable by an algebraic space locally of finite type. The $\infty$-category of commutative locally algebraic group stacks is denoted by $\mathbf{AbSt}^{lalg}_{k}\subset\mathbf{AbSt}_{k}$.
Unlike algebraic group stacks, the locally algebraic stacks are stable by suspension: if $E\in\mathbf{AbSt}_{k}^{lalg}$, then $BE$, computed in $\mathbf{AbSt}_{k}$, remains in $\mathbf{AbSt}_{k}^{lalg}$. More generally, $\mathbf{AbSt}_{k}^{lalg}$ is stable by finite limits and colimits, as this can be easily seen using the long exact sequence in homotopy.
@@ 2. Algebraic homology
defined by $E\mapsto\mathbf{Map}(X,E)$, and where $\mathsf{Top}$ denotes the $\infty$-category of Kan simplicial sets. This $\infty$-functor will be denoted by $E\mapsto\mathbb{H}(X,E)$. By definition it obviously commutes with finite limits, and thus defines a pro-object in $\mathbf{AbSt}_{k}^{alg}$ (see [18, Def. 7.1.6.1] and [18, Rem. 7.1.6.2])
@@
By definition we have in particular that for any commutative group scheme $G$ of finite type over $k$
@@
The notion above of algebraic homology is very close to that of flat homology of [20], but is different. The major difference is that these two objects live in two different, non-equivalent, $\infty$-categories. Flat homology lives in the $\infty$-category of complexes of commutative pro-algebraic groups. This has a canonical $\infty$-functor to $\mathbf{AbSt}_{k}$, by considering the stack pro-represented by a complex of pro-algebraic groups, but this $\infty$-functor is known not to be fully faithful. This is due to the existence of two different, non-isomorphic, definitions of ext-groups of commutative algebraic groups: Yoneda ext-groups inside algebraic groups and ext-groups of the corresponding fppf-sheaves. It is known that these two notions are different (see for instance [4]).
@@
Functoriality. The construction $X\mapsto\mathbb{H}^{alg}(X)$ is functorial in $X$, and defines an $\infty$-functor from Artin $k$-stacks to pro-objects in derived commutative algebraic groups. Moreover, the identity of $\mathbb{H}^{alg}(X)$ defines a canonical class $\alpha_{X}\in\mathbb{H}(X,\mathbb{H}^{alg}(X))$, and thus a canonical morphism of pro-stacks
@@
Homology groups. The pro-object $\mathbb{H}^{alg}(X)$ possesses pro-homology groups $\mathbb{H}_{i}^{alg}(X):=\pi_{i}(\mathbb{H}^{alg}(X))$, which are commutative pro-algebraic groups over $k$. By definition of our notion of commutative algebraic group stacks $\mathbb{H}_{0}^{alg}(X)$ is a pro-group scheme locally of finite presentation, and all the $\mathbb{H}_{i}^{alg}(X)$ are pro-group schemes of finite type for $i>0$.
@@
$0\to\mathbb{H}_{0}^{alg}(X)^{(0)}\to\mathbb{H}_{0}^{alg}(X)\to\mathbb{Z}\to 0.$
We can consider the universal morphism $X\to\mathbb{H}^{alg}(X)$, and consider the composite morphism $X\to\mathbb{H}_{0}^{alg}(X)\to\mathbb{Z}$. As $\mathbb{H}^{alg}(-)$ is functorial, this morphism is the composition $X\to\mathbf{Spec}\,k\to\mathbb{H}^{alg}(\mathbf{Spec}\,k)=\mathbb{Z}$. By definition the second morphism $\mathbf{Spec}\,k\to\mathbb{Z}$ exhibits $\mathbb{Z}$ as the free sheaf of abelian groups generated by $\mathbf{Spec}\,k$, and thus corresponds to the global section $1\in\mathbb{Z}(k)$. As a consequence, the composition $X\to\mathbb{Z}$ is the global function $1$ on $X$.
As a consequence of the discussion above, the morphism $X\to\mathbb{H}^{alg}(X)$ factors through a canonical morphism
@@
where $\mathbb{H}_{0}^{alg}(X)^{(1)}$ is the fiber at $1$ of the short exact sequence above, considered naturally as a torsor over the group $\mathbb{H}_{0}^{alg}(X)^{(0)}$. This morphism is a generalization of the usual Albanese morphism to the Albanese variety. When $X$ is smooth and proper, $\mathbb{H}_{0}^{alg}(X)^{(1)}$ can be identified with the usual Albanese variety, and $alb_{X}$ with the usual the Albanese map, as these two objects satisfy the same universal property. In general, the universal morphism $X\to\mathbb{H}^{alg}(X)$ can be understood as a higher cohomological generalization of the Albanese morphism as being the universal map towards a torsor over an object in $Pro(\mathbf{AbSt}_{k}^{alg})$.
Relation to the Albanese stack. When $X$ is an Artin stack of finite type over $X$, the first Postnikov truncation $\mathbb{H}_{\leq 1}^{alg}(X)$ is an abelian group object in algebraic pro-Artin $1$-stacks. This object is closely related to $D(Pic(X))$, the dual of the Picard stack of $X$ (relative to $k$) studied in [7]. Indeed, $Pic(X):=\underline{\mathbf{Map}}_{St_{k}}(X,B\mathbb{G}_{m})$ is the internal $Hom$ stack from $X$ to $B\mathbb{G}_{m}$, and we thus have a canonical evaluation morphism $ev:X\times Pic(X)\to B\mathbb{G}_{m}$ which defines a canonical morphism
$X\to\underline{\mathbf{Map}}_{Ab_{k}}(Pic(X),B\mathbb{G}_{m})=D(Pic(X)),$
where $\underline{\mathbf{Map}}_{Ab_{k}}$ denotes the stack of abelian group stacks morphisms. When $X$ is proper and the dual $D(Pic(X))$ is representable by an algebraic group stack, there is a fibration sequence of commutative group stacks
$0\to D(\pi_{0}Pic(X))\to D(Pic(X))\to\mathbb{Z}\to 0.$
This fibration sequence is obtained by mapping towards $B\mathbb{G}_{m}$ the fibration sequence
$0\to B\mathbb{G}_{m}=Pic(\mathbf{Spec}\,k)\to Pic(X)\to\pi_{0}Pic(X)\to 0.$
The choice of a global point $x\in X(k)$ defines a splitting of these two sequences, and thus the evaluation map $X\to D(Pic(X))$ can be composed to a morphism $X\to D(\pi_{0}Pic(X)$. By the universal property of algebraic homology we get a canonical morphism of pro-objects in commutative algebraic $1$-stacks
$\mathbb{H}_{\leq 1}^{alg}(X)^{(0)}\to D(\pi_{0}Pic(X)).$
In absence of global points we get a morphism $\mathbb{H}_{\leq 1}^{alg}(X)^{(1)}\to D(\pi_{0}Pic(X))^{(1)}$ where $\mathbb{H}_{\leq 1}^{alg}(X)^{(1)}$ (resp. $D(\pi_{0}Pic(X))^{(1)}$) is the fiber at $1$ of the canonical fibration sequence
@@
$\left(\textrm{resp.}\;D(\pi_{0}Pic(X))\to D(Pic(X))\to\mathbb{Z}\right).$
This map is an isomorphism when $X$ satisfies some extra conditions expressed in [7] (e.g. when $X$ is smooth and proper over $k$), as both objects $\mathbb{H}_{\leq 1}^{alg}(X)^{(1)}$ and $D(\pi_{0}Pic(X))^{(1)}$ then satisfy the same universal property. This provides a direct relation between algebraic homology and the Albanese morphism constructed in [7].
@@
We claim that this morphism is in fact an equivalence of pro-objects. Indeed, for any $G\in AbSt_{k^{\prime}}^{alg}$, we have